OpenStudy (anonymous):

So, I have a few more questions? Let 7+(2/x)+4/(x^2) Find f'(x). Find f'(5). I can't seem to grasp the concept of derivatives. I understand a little, but the problems that continue to take up a lot of space seem to confuse me. Can someone maybe walk me through this one?

OpenStudy (anonymous):

you plug in 1 for x first and then resolve problem plug in 5 for x need help solving further message me

OpenStudy (anonymous):

okay. So, should I have gotten 13, and 189/25? If not, I probably have to message you. lol

OpenStudy (anonymous):

yes u did it right

OpenStudy (anonymous):

Oh. Because in calc class, we have to use the derivitative equation, f'(x)=f(x+h)-f(x)/h so I'm a little confused?

OpenStudy (anonymous):

oh ur in calculus your smarter than me.. sorry I prob told u wrong then... maybe someone else can help you use the derivative equation?? SOrry

OpenStudy (anonymous):

or for further resources u can use www.ask.com www.wolframalpha.com or ur book or teacher tom... Sorry I couldn't help more so sorry

OpenStudy (anonymous):

haha no, you're fine. A lot of people have been telling me to "plug this in" and they've been right so haha it's all good. thanks, though!

OpenStudy (anonymous):

you have to find the derivative first then plug it in!

OpenStudy (anonymous):

is it \[f(x)=7+\frac{2}{x}+\frac{4}{x^2}\]?

OpenStudy (anonymous):

yes!

jigglypuff314 (jigglypuff314):

ah yeah, I'm in calc class right now too, we just learned the short cut yesterday :P the hint to the short cut was y=3x^100 y'=300x^99

OpenStudy (anonymous):

I'm slow and I don't understand the shortcut lol

jigglypuff314 (jigglypuff314):

\[deriv( x ^{n})=n \times x ^{n-1}\]

jigglypuff314 (jigglypuff314):

so \[f(x)=7+\frac{ 2 }{ x }+\frac{ 4 }{ x ^{2} } \rightarrow f(x)=7+2x ^{-1}+4x^{-2}\] \[f'(x)=-2x ^{-2}-8x^{-3}\]

OpenStudy (anonymous):

7 becomes 0, right?

jigglypuff314 (jigglypuff314):

yes

OpenStudy (anonymous):

Okay, yes, thank you very very much, man. Appreciate it muchly.

OpenStudy (anonymous):

for f'(5) does it equal -2/25-8/125, get like denom, them subtract.

jigglypuff314 (jigglypuff314):

I'm not sure :/ sorry

OpenStudy (anonymous):

it's cool, bro! You helped me way more than I could've helped myself! thanks again!